# Composite Flow Matching for Reinforcement Learning with Shifted-Dynamics Data

### v2

@@ Abstract
Incorporating pre-collected offline data from a source environment can significantly improve the sample efficiency of reinforcement learning (RL), but this benefit is often challenged by discrepancies between the transition dynamics of the source and target environments. Existing methods typically address this issue by penalizing or filtering out source transitions in high dynamics-gap regions. However, their estimation of the dynamics gap often relies on KL divergence or mutual information, which can be ill-defined when the source and target dynamics have disjoint support. To overcome these limitations, we propose CompFlow, a method grounded in the theoretical connection between flow matching and optimal transport. Specifically, we model the target dynamics as a conditional flow built upon the output distribution of the source-domain flow, rather than learning it directly from a Gaussian prior. This composite structure offers two key advantages: (1) improved generalization for learning target dynamics, and (2) a principled estimation of the dynamics gap via the Wasserstein distance between source and target transitions. Leveraging our principled estimation of the dynamics gap, we further introduce an optimistic active data collection strategy that prioritizes exploration in regions of high dynamics gap, and theoretically prove that it reduces the performance disparity with the optimal policy. Empirically, CompFlow outperforms strong baselines across several RL benchmarks with shifted dynamics.
@@ 1 Introduction
Reinforcement Learning (RL) has demonstrated the ability to train policies capable of achieving remarkable performance across a diverse range of complex sequential decision-making tasks [47, 49]. However, in many real-world applications, such as robotics [25, 54], healthcare [63], or wildlife conservation [60, 61, 26], the opportunity for extensive interactions with the environment is either prohibitively expensive, unsafe, or simply unavailable, thereby hindering efficient policy learning. To address this issue and enhance sample efficiency during online training, one promising approach involves merging a pre-collected, static offline dataset [53, 39]. This allows the agent to learn from a richer set of experiences, potentially accelerating convergence and improving initial performance.
@@
A critical challenge in RL with using offline data arises when the dataset is collected from source environments whose dynamics differ from those of the target domain, where the agent is simultaneously interacting and learning [44]. This issue, commonly referred to as shifted dynamics, can introduce severe mismatches that bias policy updates, destabilize the learning process, and ultimately degrade performance. For instance, in the conservation domain, historical data collected from one national park may not transfer well to another, as geography and poacher behavior can vary significantly across regions [60]. In robotics, differences in physical parameters such as friction can result in transition dynamics that diverge from the source domain.
@@
To address these challenges, existing methods either penalize the rewards or value estimates of source transitions with high dynamics gap [41, 32], or filter out such transitions entirely [58]. However, these approaches face key limitations. Most notably, the estimation of the dynamics gap typically relies on KL divergence or mutual information, both of which can be ill-defined when the source and target transition dynamics have disjoint support [2, 43].
@@
Figure 1: Comparison between direct and composite flow matching. Composite flow first transports from a Gaussian latent variable to the source domain distribution, then adapts to the target distribution via optimal transport flow matching.
@@
In this paper, we propose CompFlow, a novel method for RL with shifted-dynamics data by leveraging the theoretical connection between flow matching and optimal transport. Specifically, we model the transition dynamics of the target domain using a composite flow architecture, where the target flow is constructed on top of the output distribution of the source flow, rather than being learned from a Gaussian prior. This design enables a principled estimation of the dynamics gap via the Wasserstein distance between the source and target transition distributions. On the theoretical side, we show that the composite flow formulation reduces generalization error by reusing structural knowledge from the source domain, particularly when target domain data is scarce—outperforming standard approaches that learn the target flow from scratch.
@@
Building on the dynamics gap estimated by composite flow matching using the Wasserstein distance, we go beyond selectively merging source transitions with low dynamics gap and further propose an active data collection strategy that targets regions exhibiting high dynamics gap relative to the source domain. Such regions are often underrepresented in the replay buffer due to the dominance of low-gap samples. Our theoretical analysis shows that targeted exploration in these high-gap regions can further close the performance gap with respect to the optimal policy.
@@
Our contributions are summarized as follows: (1) We introduce a composite flow model that estimates the dynamics gap by computing the Wasserstein distance between conditional transition distributions. We provide theoretical analysis showing that this approach achieves lower generalization error compared to learning the target dynamics from scratch. (2) Leveraging this principled estimation, we propose a new data collection strategy that encourages the policy to actively explore regions with high dynamics gap. We also provide a theoretical analysis of its performance benefits. (3) We empirically validate our method on various public RL benchmarks with shifted dynamics and demonstrate that CompFlow outperforms or matches state-of-the-art baselines across these tasks.
@@ 2.1 Problem Definition
We consider two infinite-horizon MDPs: $\mathcal{M}_{\text{src}}:=(\mathcal{S},\mathcal{A},p_{\text{src}},r,\gamma)$ and $\mathcal{M}_{\text{tar}}:=(\mathcal{S},\mathcal{A},p_{\text{tar}},r,\gamma)$, sharing the same state/action spaces, reward function $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$, and discount factor $\gamma\in(0,1)$, but differing in transition dynamics:
$p_{\text{src}}(s^{\prime}|s,a)\neq p_{\text{tar}}(s^{\prime}|s,a),\quad\text{for some }(s,a).$
We assume that the reward signals are bounded, i.e., $|r(s,a)|\leq r_{\rm max},\forall s,a$. For any policy $\pi$, we define its discounted state-action visitation under MDP $\mathcal{M}$ as $\rho_{\mathcal{M}}^{\pi}(s,a):=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}P_{\mathcal{M},t}^{\pi}(s)\pi(a\mid s)$, and the corresponding state visitation as $d_{\mathcal{M}}^{\pi}(s):=\sum_{a}\rho_{\mathcal{M}}^{\pi}(s,a)$. The expected return is $\eta_{\mathcal{M}}(\pi):=\mathbb{E}_{(s,a)\sim\rho_{\mathcal{M}}^{\pi}}[r(s,a)]$.
Definition 2.1 (Online Dynamics Adaptation with Offline Data).
Given an offline dataset $\mathcal{D}_{\text{src}}=\{(s_{i},a_{i},s^{\prime}_{i},r_{i})\}_{i=1}^{N}$ from $\mathcal{M}_{\text{src}}$ and limited online access to $\mathcal{M}_{\text{tar}}$, the objective is to learn a policy $\pi$ maximizing $\eta_{\mathcal{M}_{\text{tar}}}(\pi)$, ideally approaching $\eta_{\mathcal{M}_{\text{tar}}}(\pi^{\star})$, where $\pi^{\star}:=\arg\max_{\pi}\eta_{\mathcal{M}_{\text{tar}}}(\pi)$.
Lemma 2.2 (Return Bound between Source Domain and Target Domain [32]).
Let the empirical behavior policy in $\mathcal{D}_{\text{src}}$ be $\pi_{\mathcal{D}_{\text{src}}}(a\mid s)$. Define $C_{1}=\frac{2r_{\max}}{(1-\gamma)^{2}}$. Then for any policy $\pi$,
$\eta_{\mathcal{M}_{\text{tar}}}(\pi)-\eta_{\mathcal{M}_{\text{src}}}(\pi)\geq\;$ $-2C_{1}\,\mathbb{E}_{(s,a)\sim\rho_{\mathcal{M}}^{\pi_{\mathcal{D}_{\text{src}}}},s^{\prime}\sim p_{\text{src}}}\left[D_{\mathrm{TV}}(\pi(\cdot|s^{\prime})\parallel\pi_{\mathcal{D}_{\text{src}}}(\cdot|s^{\prime}))\right]$
$-C_{1}\,\mathbb{E}_{(s,a)\sim\rho_{\mathcal{M}}^{\pi_{\mathcal{D}_{\text{src}}}}}\left[D_{\mathrm{TV}}(p_{\text{tar}}(\cdot|s,a)\parallel p_{\text{src}}(\cdot|s,a))\right].$
This bound highlights two key sources of return gap between domains: (1) the mismatch between the learned policy and the behavior policy in the offline dataset, and (2) the shift in environment dynamics. The former can be mitigated through behavior cloning [32, 59], while the latter can be addressed by filtering out source transitions with large dynamics gaps [58, 34]. A central challenge lies in accurately estimating this gap. Existing methods often rely on KL divergence or mutual information [41, 32], which can be ill-defined when the two dynamics have disjoint support.
@@ 2.2 Flow Matching
In this paper, we will adopt flow matching [29, 31, 1] to model transition dynamics in reinforcement learning, owing to its ability to capture complex distributions. Flow Matching (FM) offers a simpler alternative to denoising diffusion models [20, 51], which are typically formulated using stochastic differential equations (SDEs). In contrast, FM is based on deterministic ordinary differential equations (ODEs), providing advantages such faster inference, and often improved sample quality.
@@
where $x_{t}=(1-t)x_{0}+tx_{1}$ denotes the linear interpolation between $x_{0}$ and $x_{1}$. Using linear interpolation paths encourages the learned flow to follow nearly straight-line trajectories, which reduces discretization error and improves the computational efficiency of ODE solvers during sampling.
@@ 2.3 Optimal Transport Flow Matching and Wasserstein Distance
For theoretical justification, see Theorem 4.2 in [42].
In practice, the optimal coupling $\pi^{\star}$ is approximated using mini-batches by solving a discrete OT problem between empirical samples from $p_{0}$ and $p_{1}$. The use of mini-batch OT has also been shown to implicitly regularize the transport plan [13, 14], as the stochasticity from independently sampled batches induces behavior similar to entropic regularization [9]. Further details of the OT-FM training procedure are provided in Appendix B.
Figure 2: Overall Framework of CompFlow. To estimate the dynamics gap, we propose composite flow matching, which computes the Wasserstein distance between source and target conditional transition distributions. Guided by the estimated dynamics gap, we augment policy training with source-domain transitions that exhibit low discrepancy from the target dynamics, and incorporate a behavior cloning objective to stabilize learning. To enhance data diversity and facilitate adaptation, we further encourage exploration in regions with high dynamics gap.
@@ 3.1 Estimating Dynamics Gap via Composite Flow
To estimate the dynamics gap, we begin by learning the transition dynamics in both the source and target domains. While flow matching offers a powerful framework for modeling complex transition dynamics, the limited number of training samples in the target domain poses a significant challenge. Naively training separate flow models for each domain can lead to poor generalization in the target domain due to data scarcity.
To mitigate this, we propose a composite flow formulation. Instead of learning the target transition model $p_{\text{tar}}(s^{\prime}|s,a)$ from scratch, we leverage structural knowledge from a well-trained source model $p_{\text{src}}(s^{\prime}|s,a)$. This enables knowledge transfer across domains while adapting to the specific characteristics of the target environment.
Source Flow. We begin by learning a conditional flow model for the source domain. Let $x_{0}\sim\mathcal{N}(0,\mathbf{I})$ be the initial latent variable. The source flow map $\psi^{\text{src}}_{\theta}(x_{0},t|s,a)$ is defined as the solution to the following ODE:
$\frac{\mathrm{d}}{\mathrm{d}t}\psi^{\text{src}}_{\theta}(x_{0},t|s,a)=v^{\text{src}}_{\theta}(\psi^{\text{src}}_{\theta}(x_{0},t|s,a),t,s,a),\quad\psi^{\text{src}}_{\theta}(x_{0},0|s,a)=x_{0}.$
Solving this ODE from $t=0$ to $1$ produces an intermediate representation: $x_{1}=\psi^{\text{src}}_{\theta}(x_{0},1|s,a).$
Target Flow. Instead of learning a target flow directly from a Gaussian prior, we initialize it from the intermediate representation $x_{1}$ produced by the source flow. The target flow map $\psi^{\text{tar}}_{\phi}(x,t|s,a)$ is defined as:
$\frac{\mathrm{d}}{\mathrm{d}t}\psi^{\text{tar}}_{\phi}(x,t|s,a)=v^{\text{tar}}_{\phi}(\psi^{\text{tar}}_{\phi}(x,t|s,a),t,s,a),\quad\psi^{\text{tar}}_{\phi}(x,1|s,a)=x_{1}.$
Solving this ODE from $t=1$ to $2$ yields the final target-domain prediction: $s^{\prime}\doteq x_{2}=\psi^{\text{tar}}_{\phi}(x_{1},2|s,a).$
@@
We assume the state space is bounded for both source and target domains, i.e., $|s|\leq\zeta\ \forall s\in\mathcal{S}$. For an unseen sample pair $(s,a)$, there exists a constant $C>0$ (independent of sample sizes and problem dimension) such that, the mean square error of the composite flow $Err_{\text{cmpst}}$ is smaller than that of the direct flow $Err_{\text{direct}}$, i.e., $Err_{\text{cmpst}}<Err_{\text{direct}}$ whenever the total-variation distance obeys the explicit bound
$D_{TV}(p_{\text{src}}(\cdot|s,a),p_{\text{tar}}(\cdot|s,a))\;\leq\;\frac{C}{\zeta^{2}}\Bigl{(}-\frac{\operatorname{Tr}(\Sigma_{s^{\prime}}^{\text{src}})}{\sqrt{N_{\text{src}}}}+\frac{\operatorname{Tr}(\Sigma_{s^{\prime}}^{\text{tar}})}{\sqrt{N_{\text{tar}}}}\Bigr{)}\ \forall s,a,$
where $\Sigma_{s^{\prime}}^{\text{tar}}$, $\Sigma_{s^{\prime}}^{\text{src}}$ are the covariance matrices of the $N_{\text{tar}}$ samples collected in the target MDP and the $N_{\text{src}}$ samples collected in the source MDP, respectively.
As shown in Theorem 3.1, if the distance between the source and target dynamics is below a certain threshold, our composite flow model achieves a smaller generalization error compared to directly learning the target flow from a Gaussian prior. This aligns with our setting: when the two dynamics deviates too large, leveraging source-domain data offers little benefit. Moreover, when the amount of target-domain data is limited (i.e., small $N_{\rm tar}$), the tolerance for this distance becomes more relaxed—consistent with our setting where interaction with the target environment is scarce. The bound also becomes more relaxed with more source domain data (larger $N_{\rm src}$), which aligns with our setting of having abundant source domain data.
To compute the Wasserstein distance between the source and target transition distributions, we can use the OT-FM objective to train the target flow, initialized from the source flow distribution. However, when the state-action pair $(s,a)$ lies in a continuous space, it is not feasible to obtain a batch of samples corresponding to one $(s,a)$ from the target domain.
To address this, we follow prior works [22, 17] that handle samples with distinct conditioning variables. The key idea is to incorporate the conditioning variables into the cost function when solving for the optimal transport coupling. Specifically, we define the cost between two sample pairs $(s_{\rm src},a_{\rm src},s^{\prime}_{\rm src})$ and $(s_{\rm tar},a_{\rm tar},s^{\prime}_{\rm tar})$ as:
$c\big{(}(s_{\rm src},a_{\rm src},s^{\prime}_{\rm src},s_{\rm tar},a_{\rm tar},s^{\prime}_{\rm tar}\big{)}=\|s^{\prime}_{\rm src}-s^{\prime}_{\rm tar}\|_{2}^{2}+\eta\left(\|s_{\rm src}-s_{\rm tar}\|_{2}^{2}+\|a_{\rm src}-a_{\rm tar}\|_{2}^{2}\right),$
@@
Then the training objective of the target flow follows
$\mathcal{L}_{\rm tar}(\phi)=\mathbb{E}_{((s_{\rm src},a_{\rm src},s^{\prime}_{\rm src}\doteq x_{1}),(s_{\rm tar},a_{\rm tar},s^{\prime}_{\rm tar}\doteq x_{2}))\sim q^{*},t\sim\mathcal{U}[1,2]}\left[\left\|v_{\phi}\left(x_{t},t,s_{t},a_{t}\right)-(s_{\rm tar}-s_{\rm src})\right\|^{2}\right],$
where $q^{*}$ is the optimal coupling of empirical distribution $(s_{\rm src},a_{\rm src},s^{\prime}_{\rm src})$ generated by learned source flow and empirical transition $(s_{\rm tar},a_{\rm tar},s^{\prime}_{\rm tar})$ from target domain replay buffer. $x_{t}$ is the linear interpolation between $x_{1}$ and $x_{2}$, $s_{t}$ is the interpolation between $s_{\rm src}$ and $s_{\rm tar}$, and $a_{t}$ is the interpolation between $a_{\rm src}$ and $a_{\rm tar}$.
The complete training algorithms of the source flow and the target flow are in Appendix B.
@@
[…]
$W_{2}^{2}(p_{\rm src}(\cdot|s,a),p_{\rm tar}(\cdot|s,a))=\mathbb{E}_{x_{0}\sim\mathcal{N}(0,\mathbf{I})}\left[\left\|\psi^{\text{src}}_{\theta}(x_{0},1|s,a)-\psi^{\text{tar}}_{\phi}(\psi^{\text{src}}_{\theta}(x_{0},1|s,a),2|s,a)\right\|_{2}^{2}\right].$
@@
$\widehat{\Delta}(s,a)=\left(\frac{1}{M}\sum_{j=1}^{M}\left\|\psi^{\text{src}}_{\theta}(x_{0}^{(j)},1|s,a)-\psi^{\text{tar}}_{\phi}(\psi^{\text{src}}_{\theta}(x_{0}^{(j)},1|s,a),2|s,a)\right\|^{2}_{2}\right)^{\frac{1}{2}},\quad x_{0}^{(j)}\sim\mathcal{N}(0,\mathbf{I}).$ (2)

### v3

@@ Abstract
Incorporating pre-collected offline data can significantly improve the sample efficiency of reinforcement learning (RL), but this benefit is often challenged by discrepancies between the transition dynamics of the offline data and the online environment. Existing methods typically address this issue by penalizing or filtering out offline transitions in high dynamics-gap regions. However, their estimation of the dynamics gap often relies on KL divergence or mutual information, which can be ill-defined when the source and target dynamics have different support. To overcome these limitations, we propose CompFlow, a method grounded in the theoretical connection between flow matching and optimal transport. Specifically, we model the online dynamics as a conditional flow built upon the output distribution of the offline flow, rather than learning it directly from a Gaussian prior. This composite structure offers two key advantages: (1) improved generalization for learning online dynamics, and (2) a principled estimation of the dynamics gap via the Wasserstein distance between offline and online transitions. Leveraging our principled estimation of the dynamics gap, we further introduce an optimistic active data collection strategy that prioritizes exploration in regions of high dynamics gap, and theoretically prove that it reduces the performance disparity with the optimal policy. Empirically, CompFlow outperforms strong baselines across various RL benchmarks with shifted dynamics.
@@ 1 Introduction
Reinforcement Learning (RL) has demonstrated remarkable performance in complex sequential decision-making tasks such as playing Go and Atari games [schrittwieser2020mastering, silver2017mastering], supported by access to large amounts of online interactions with the environment. However, in many real-world domains such as robotics [kober2013reinforcement, tang2025deep], healthcare [yu2021reinforcement], and wildlife conservation [xu2021dual, xu2021robust, kong2025robust, kong2025generative], access to such interactions is often prohibitively expensive, unsafe, or infeasible. The limited availability of interactions presents a major challenge for learning effective and reliable policies. To address this challenge and improve sample efficiency during online training, a promising strategy is to incorporate a pre-collected offline dataset generated by a previous policy [song2022hybrid, nakamoto2023cal]. This approach enables the agent to learn from a broader set of experiences, which can help accelerate learning and improve performance [song2022hybrid].
@@
A critical challenge in RL with offline data arises when the transition dynamics of the offline dataset differ from those of the online environment, where the agent is actively interacting and learning [qu2025hybrid]. This issue, commonly referred to as shifted dynamics, can introduce severe mismatches that bias policy updates, destabilize the learning process, and ultimately degrade performance. For instance,In robotics, differences in physical parameters such as friction can result in transition dynamics that diverge from the source domain. For example, in the conservation domain, historical data collected from one region may not generalize well to another, as differences in environmental conditions and ecological patterns across regions can lead to variations in poaching behavior [xu2021dual].
@@
To address these challenges, existing methods either penalize the rewards or value estimates of offline transitions with high dynamics gap [niu2022trust, lyu2024cross], or filter out such transitions entirely [wen2024contrastive]. However, these approaches face key limitations. Most notably, the estimation of the dynamics gap typically relies on KL divergence or mutual information, both of which can be ill-defined when the offline and online transition dynamics have different supports [arjovsky2017wasserstein, poole2019variational].
@@
Figure 1: Comparison between direct and composite flow matching. Composite flow first transports from a Gaussian latent variable to the offline transition distribution, then adapts to the online distribution via optimal transport flow matching.
@@
In this paper, we propose CompFlow, a new method for RL with shifted-dynamics data that leverages the theoretical connection between flow matching and optimal transport. We model the transition dynamics of the online environment using a composite flow architecture, where the online flow is defined on top of the output distribution of a learned offline flow rather than being initialized from a Gaussian prior. This design enables a principled estimation of the dynamics gap using the Wasserstein distance between the offline and online transition dynamics. On the theoretical side, we show that the composite flow formulation reduces generalization error compared standard flow matching by reusing structural knowledge embedded in the offline data, particularly when online interactions are limited.
@@
Building on the dynamics gap estimated by composite flow matching using the Wasserstein distance, we go beyond selectively merging offline transitions with low dynamics gap and further propose an active data collection strategy that targets regions in the online environment where the dynamics gap relative to the offline data is high. Such regions are often underrepresented in the replay buffer due to the dominance of low-gap samples. Our theoretical analysis shows that targeted exploration in these high-gap regions can further close the performance gap with respect to the optimal policy.
@@
Our contributions are summarized as follows: (1) We introduce a composite flow model that estimates the dynamics gap by computing the Wasserstein distance between conditional transition distributions. We provide theoretical analysis showing that this approach achieves lower generalization error compared to learning the online dynamics from scratch. (2) Leveraging this principled estimation, we propose a new data collection strategy that encourages the policy to actively explore regions with high dynamics gap. We also provide a theoretical analysis of its performance benefits. (3) We empirically validate our method on various RL benchmarks with shifted dynamics and demonstrate that CompFlow outperforms or matches state-of-the-art baselines across these tasks.
@@ 2.1 Problem Definition
We consider two infinite-horizon MDPs: $\mathcal{M}_{\text{off}}:=(\mathcal{S},\mathcal{A},p_{\text{off}},r,\gamma)$ and $\mathcal{M}_{\text{on}}:=(\mathcal{S},\mathcal{A},p_{\text{on}},r,\gamma)$, sharing the same state/action spaces, reward function $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$, and discount factor $\gamma\in(0,1)$, but differing in transition dynamics:
$p_{\text{off}}(s^{\prime}|s,a)\neq p_{\text{on}}(s^{\prime}|s,a)\quad\text{for some }(s,a).$
We assume rewards are bounded, i.e., $|r(s,a)|\leq r_{\max}$ for all $s,a$. For any policy $\pi$, let $(s_{t},a_{t})_{t\geq 0}$ be the trajectory generated by $\mathcal{M}$ and $\pi$. We define the discounted state–action visitation (occupancy) measure as $\rho_{\mathcal{M}}^{\pi}(s,a):=(1-\gamma)\,\mathbb{E}\!\left[\sum_{t=0}^{\infty}\gamma^{t}\,\mathbf{1}\{s_{t}=s,\,a_{t}=a\}\right]$, and the discounted state visitation as $d_{\mathcal{M}}^{\pi}(s):=\sum_{a}\rho_{\mathcal{M}}^{\pi}(s,a).$ The expected return is $\eta_{\mathcal{M}}(\pi):=\mathbb{E}_{(s,a)\sim\rho_{\mathcal{M}}^{\pi}}[r(s,a)]$.
Definition 2.1 (Online Policy Learning with Shifted-Dynamics Offline Data).
Given an offline dataset $\mathcal{D}_{\text{off}}=\{(s_{i},a_{i},s^{\prime}_{i},r_{i})\}_{i=1}^{N}$ from $\mathcal{M}_{\text{off}}$ and limited online access to $\mathcal{M}_{\text{on}}$, the objective is to learn a policy $\pi$ maximizing $\eta_{\mathcal{M}_{\text{on}}}(\pi)$, ideally approaching $\eta_{\mathcal{M}_{\text{on}}}(\pi^{\star})$, where $\pi^{\star}:=\arg\max_{\pi}\eta_{\mathcal{M}_{\text{on}}}(\pi)$.
Lemma 2.2 (Return Bound between Two Environments [lyu2024cross]).
Let the empirical behavior policy in $\mathcal{D}_{\rm off}$ be $\pi_{\mathcal{D}_{\rm off}}(a\mid s)$. Define $C_{1}=\frac{2r_{\max}}{(1-\gamma)^{2}}$. Then for any policy $\pi$,
$\eta_{\mathcal{M}_{\rm on}}(\pi)-\eta_{\mathcal{M}_{\rm off}}(\pi)\geq\;$ $-2C_{1}\,\mathbb{E}_{(s,a)\sim\rho_{\mathcal{M}}^{\pi_{\mathcal{D}_{\rm off}}},s^{\prime}\sim p_{\rm off}}\left[D_{\mathrm{TV}}(\pi(\cdot|s^{\prime})\parallel\pi_{\mathcal{D}_{\rm off}}(\cdot|s^{\prime}))\right]$
$-C_{1}\,\mathbb{E}_{(s,a)\sim\rho_{\mathcal{M}}^{\pi_{\mathcal{D}_{\rm off}}}}\left[D_{\mathrm{TV}}(p_{\rm on}(\cdot|s,a)\parallel p_{\rm off}(\cdot|s,a))\right].$
This bound highlights two key sources of return gap between domains: (1) the mismatch between the learned policy and the behavior policy in the offline dataset, and (2) the shift in environment dynamics. The former can be mitigated through behavior cloning [lyu2024cross, xu2023cross], while the latter can be addressed by filtering out source transitions with large dynamics gaps [wen2024contrastive, lyu2025cross]. A central challenge lies in accurately estimating this gap. Existing methods often rely on KL divergence or mutual information [niu2022trust, lyu2024cross], which can be ill-defined when the two dynamics have different supports.
@@ 2.2 Flow Matching
In this paper, we will adopt flow matching [lipman2023flow, liu2022flow, albergo2023building] to model transition dynamics in reinforcement learning, owing to its ability to capture complex distributions. Flow Matching (FM) offers a simpler alternative to denoising diffusion models [ho2020denoising, song2020score], which are typically formulated using stochastic differential equations (SDEs). In contrast, FM is based on deterministic ordinary differential equations (ODEs), providing advantages such faster inference, and often improved sample quality.
@@
where $x_{t}=(1-t)x_{0}+tx_{1}$ denotes the linear interpolation between $x_{0}$ and $x_{1}$. Using linear interpolation paths encourages the learned flow to follow nearly straight-line trajectories, which reduces discretization error and improves the computational efficiency of ODE solvers during sampling [liu2022flow].
@@ 2.3 Optimal Transport Flow Matching and Wasserstein Distance
For theoretical justification, see Theorem 4.2 in [pooladian2023multisample].
In practice, the optimal coupling $\pi^{\star}$ is approximated using mini-batches by solving a discrete OT problem between empirical samples from $p_{0}$ and $p_{1}$. The use of mini-batch OT has also been shown to implicitly regularize the transport plan [fatras2019learning, fatras2021minibatch], as the stochasticity from independently sampled batches induces behavior similar to entropic regularization [cuturi2013sinkhorn]. Further details of the OT-FM training procedure are provided in Appendix C.
Figure 2: Overall Framework of CompFlow. To estimate the dynamics gap, we propose composite flow matching, which computes the Wasserstein distance between offline and online transition dynamics. Guided by the estimated dynamics gap, we augment policy training with offline transitions that exhibit low discrepancy from the online dynamics, and incorporate a behavior cloning objective to stabilize learning. To enhance data diversity and facilitate adaptation, we further encourage exploration in regions with high dynamics gap.
@@ 3.1 Estimating Dynamics Gap via Composite Flow
To estimate the gap between dynamics, we first learn the transition models for both the offline dataset and the online environment. Flow matching provides a flexible framework for modeling complex transition dynamics; however, the limited number of samples available in the online environment presents a significant challenge. Training separate flow models for each environment can result in poor generalization in the online environment, as the model may overfit to the small amount of available data.
To mitigate this, we propose a composite flow formulation. Instead of learning the online transition model $p_{\text{on}}(s^{\prime}|s,a)$ from scratch, we leverage structural knowledge from a well-trained offline model $p_{\text{off}}(s^{\prime}|s,a)$. This enables to incorporate prior knowledge to improve the generalization.
Offline flow. We begin by learning a conditional flow model for the offline data. Let $x_{0}\sim\mathcal{N}(0,\mathbf{I})$ be the initial latent variable. The offline flow map $\psi^{\text{off}}_{\theta}(x_{0},t|s,a)$ is defined as the solution to the following ODE:
$\frac{\mathrm{d}}{\mathrm{d}t}\psi^{\text{off}}_{\theta}(x_{0},t|s,a)=v^{\text{off}}_{\theta}(\psi^{\text{off}}_{\theta}(x_{0},t|s,a),t,s,a),\quad\psi^{\text{off}}_{\theta}(x_{0},0|s,a)=x_{0}.$
Solving this ODE from $t=0$ to $1$ produces an intermediate representation: $x_{1}=\psi^{\text{off}}_{\theta}(x_{0},1|s,a).$
Online flow. Instead of learning a online flow directly from a Gaussian prior, we initialize it from the intermediate representation $x_{1}$ produced by the offline flow. The target flow map $\psi^{\text{on}}_{\phi}(x,t|s,a)$ is defined as:
$\frac{\mathrm{d}}{\mathrm{d}t}\psi^{\text{on}}_{\phi}(x,t|s,a)=v^{\text{on}}_{\phi}(\psi^{\text{on}}_{\phi}(x,t|s,a),t,s,a),\quad\psi^{\text{on}}_{\phi}(x,1|s,a)=x_{1}.$
Solving this ODE from $t=1$ to $2$ yields the final prediction for the online environment: $s^{\prime}\doteq x_{2}=\psi^{\text{on}}_{\phi}(x_{1},2|s,a).$
We now show that when the offline flow induces a distribution $\hat{p}_{\mathrm{off}}(s^{\prime}|s,a)$ that is closer to $p_{\mathrm{on}}(s^{\prime}|s,a)$ than the standard Gaussian distribution, the proposed composite flow method enjoys a smaller generalization error bound.
@@
Assume that composite flow and direct flow share the same hypothesis class $\mathcal{H}$ of (measurable) vector fields $v:\mathbb{R}^{d}\times[0,1]\to\mathbb{R}^{d}$, and that there exists $B\in(0,\infty)$ such that $\sup_{v\in\mathcal{H}}\sup_{x\in\mathbb{R}^{d},\,t\in[0,1]}\left\lVert v(x,t)\right\rVert_{2}\leq B$. Also, assume that $\max\{\mathrm{Tr}(\Sigma_{\rm on}),\mathrm{Tr}(\widehat{\Sigma}_{\rm off})\}\leq C_{\rm TR}$ for some $C_{\rm TR}$. The composite flow enjoys a strictly tighter high-probability generalization bound than the direct flow if and only if
$W_{2}(p_{G},p_{\rm on})>W_{2}(\hat{p}_{\rm off},p_{\rm on})$ (2)
Here, $p_{G}:=\mathcal{N}(0,I_{d})$, $\Sigma_{\rm on}$ is the covariance of $p_{\rm on}$, and $\widehat{\Sigma}_{\rm off}$ is the covariance of $\hat{p}_{\mathrm{off}}$.
Remark 3.2.
In our setting, we assume that the offline data and the online environment share meaningful similarities. Given the abundance of offline data available to accurately learn the offline flow, this assumption is expected to hold well.
To compute the Wasserstein distance between the offline and online transition dynamics, we can use the OT-FM objective to train the online flow, initialized from the offline flow distribution. However, when the state-action pair $(s,a)$ lies in a continuous space, it is not feasible to obtain a batch of samples corresponding to a fixed $(s,a)$ from the online environment.
To address this, we follow prior works [kerrigan2024dynamic, generale2024conditional] that handle samples with distinct conditioning variables. The key idea is to incorporate the conditioning variables into the cost function when solving for the optimal transport coupling. Specifically, we define the cost between two sample pairs $(s_{\rm off},a_{\rm off},s^{\prime}_{\rm off})$ and $(s_{\rm on},a_{\rm on},s^{\prime}_{\rm on})$ as:
$c\big((s_{\rm off},a_{\rm off},s^{\prime}_{\rm off},s_{\rm on},a_{\rm on},s^{\prime}_{\rm on}\big)=\|s^{\prime}_{\rm off}-s^{\prime}_{\rm on}\|_{2}^{2}+\eta\left(\|s_{\rm off}-s_{\rm on}\|_{2}^{2}+\|a_{\rm off}-a_{\rm on}\|_{2}^{2}\right),$
@@
Then the training objective of the online flow follows
$\mathcal{L}_{\rm on}(\phi)=\mathbb{E}_{((s_{\rm off},a_{\rm off},s^{\prime}_{\rm off}\doteq x_{1}),(s_{\rm on},a_{\rm on},s^{\prime}_{\rm on}\doteq x_{2}))\sim q^{*},t\sim\mathcal{U}[1,2]}\left[\left\|v_{\phi}\left(x_{t},t,s_{\rm on},a_{\rm on}\right)-(s^{\prime}_{\rm on}-s^{\prime}_{\rm off})\right\|^{2}\right],$
where $q^{*}$ is the optimal coupling of empirical distribution $(s_{\rm off},a_{\rm off},s^{\prime}_{\rm off})$ generated by learned source flow and empirical transition $(s_{\rm on},a_{\rm on},s^{\prime}_{\rm on})$ from target domain replay buffer. $x_{t}$ is the linear interpolation between $x_{1}$ and $x_{2}$.
The complete training algorithms of the offline flow and the online flow are in Appendix C.
@@
Given $(s,a)\in\mathcal{S}\times\mathcal{A}$. As $\eta\to\infty$ and the training batch size $\to\infty$, the Wasserstein distance between the source and target transition distributions satisfies
$W_{2}^{2}(p_{\rm off}(\cdot|s,a),p_{\rm on}(\cdot|s,a))=\mathbb{E}_{x_{0}\sim\mathcal{N}(0,\mathbf{I})}\left[\left\|\psi^{\text{off}}_{\theta}(x_{0},1|s,a)-\psi^{\text{on}}_{\phi}(\psi^{\text{off}}_{\theta}(x_{0},1|s,a),2|s,a)\right\|_{2}^{2}\right].$
@@
$\widehat{\Delta}(s,a)=\left(\frac{1}{M}\sum_{j=1}^{M}\left\|\psi^{\text{off}}_{\theta}(x_{0}^{(j)},1|s,a)-\psi^{\text{on}}_{\phi}(\psi^{\text{off}}_{\theta}(x_{0}^{(j)},1|s,a),2|s,a)\right\|^{2}_{2}\right)^{\frac{1}{2}},\quad x_{0}^{(j)}\sim\mathcal{N}(0,\mathbf{I}).$ (3)